in a multiple choice test each question has four options. students will get 5 points for each correct answer; lose 1 point for each incorrect answer; and recieve no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show calculations for expected value.
Answer by stanbon(57356)

in a multiple choice test each question has four options. students will get 5 points for each correct answer; lose 1 point for each incorrect answer; and recieve no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show calculations for expected value.
----------
E(points) = (1/4)5 + (3/4)(-1) = (5/4) - (3/4) = 1/2
--
Expect to gain 1/2 point on each question answered.
------------
Cheers,
Stan H.